OpenStudy (anonymous):

Write an equation for the given function given the period, phase shift, and vertical shift.

OpenStudy (anonymous):

cotangent function, period = 1/3pi , phase shift = 1/8pi , vertical shift = –9

OpenStudy (anonymous):

F(x)=1/3pi cot (x-1/8pi)-9